OpenStudy (lovelyharmonics):

When 190 grams of titanium tetrachloride react with an excess of water, as shown in the unbalanced chemical equation below, how many moles of hydrochloric acid will be produced? TiCl4 + H2O TiO2 + HCl

OpenStudy (abmon98):

TiCl4+2H2O-->TiO2+4HCl

OpenStudy (abmon98):

Mass Given:190g Mr of TiCl4:48+(35.5*4)=190(g/mol) Number of Moles=Mass/Molar Mass 190/190=1mol ratio of mole of TiCl4 to HCl is 1:4 so if 1 mole of TiCl4 reacts so 1*4/1 moles of HCl is produced

OpenStudy (lovelyharmonics):

So 0.25 moles would be produced c:

OpenStudy (abmon98):

nope 4 moles